[Hallc_running] [New Logentry] RC report - 7/16/2022

arunts at jlab.org arunts at jlab.org
Sat Jul 16 13:40:02 EDT 2022

Logentry Text:
We finished taking dummy runs, and continued to take production data on LH2 during day and swing shift yesterday. After many problems with fluctuating SHMS Q2 LN2 level which started around 00:30 (https://logbooks.jlab.org/entry/4016207)), Steve Lassiter asked us to temporarily silence the alarms and continue data taking. He power cycled the level meter this morning after the magnet was ramped down and that seemed to bring the LN2 Q2 levels back into range. At around 8:00 AM, LN2 Q1 pressure started spiking as a consequence of the Q2 levels. This has also stabilized after the level meter power cycling and is currently in range (https://logbooks.jlab.org/entry/4016280). We continued to take production data afterwards and ended run 15554 to move on to the next kinematic setting (SHMS left --9.65 deg). Steve Lassiter remotely enabled SHMS rotation and we successfully finished our spectrometer rotation. MCC will start with 60 uA at this setting and determine (and will continue t!
 o provide us) the most stable beam current. For the next few days, the plan is to continue taking production data on LH2 with an occasional dummy run and a 3/4 HMS run. 


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4016350
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220716/8acef1a6/attachment.html>

More information about the Hallc_running mailing list